3. The acoustic impedance of a certain soft tissue is 1.63 x 106 kgm-2s-1 and its density is 1.09 x 103 kgm-3. Calculate the velocity at which the ultrasound waves travel in this medium


1
Expert's answer
2021-03-23T11:11:53-0400
"Z=\rho v\\1.63 \cdot 10^6=1.09\cdot 10^3 v\\v=1500\frac{m}{s}"